# Collage: Decomposable Rapid Prototyping for Information Extraction on Scientific PDFs

### v1

@@ Abstract
Recent years in NLP have seen the continued development of domain-specific information extraction tools for scientific documents, alongside the release of increasingly multimodal pretrained transformer models. While the opportunity for scientists outside of NLP to evaluate and apply such systems to their own domains has never been clearer, these models are difficult to compare: they accept different input formats, are often black-box and give little insight into processing failures, and rarely handle PDF documents, the most common format of scientific publication. In this work, we present Collage, a tool designed for rapid prototyping, visualization, and evaluation of different information extraction models on scientific PDFs. Collage allows the use and evaluation of any HuggingFace token classifier, several LLMs, and multiple other task-specific models out of the box, and provides extensible software interfaces to accelerate experimentation with new models. Further, we enable both developers and users of NLP-based tools to inspect, debug, and better understand modeling pipelines by providing granular views of intermediate states of processing. We demonstrate our system in the context of information extraction to assist with literature review in materials science.
Collage: Decomposable Rapid Prototyping for Information Extraction on Scientific PDFs
@@ 1 Introduction
This paper presents Collage, a tool that facilitates the rapid prototyping, comparison, and evaluation of multiple models (whether text-based or multimodal) on the contents of scientific PDF documents. Collage was designed to address the interface between developers of NLP-based tools and the users of those tools. To address user needs, we ground our design in a series of interviews with domain experts in multiple fields, with a particular focus on materials science. Further, in cases where model results may not meet users’ or developers’ expectations, we visualize the intermediate representation at each step, giving the user a debuggable view of the modeling pipeline, allowing shared debugging processes between developers and users. Collage is domain-agnostic, and can visualize any model that conforms to one of its three interfaces - for token classification models, text generation models, and image/text multimodal models. We provide implementations of these interfaces that allow the use of any HuggingFace token classifier, multiple LLMs, and several additional models without requiring users to write any code. All of the interfaces are easily implemented, and we provide instructions in our repository.
We make the code available on GitHub, the demo video on YouTube, and present a running instance on our server here. If the server is down, our demo can be run through Docker Compose, with instructions in our repository.
@@ 2 Motivation
One of the most prominent themes to emerge in our interviews is the variety of focuses that scientists, even in very closely related subfields, can have when reading a paper and evaluating it for relevance to their purpose. While many participants focused on paper metadata, such as the reputation of the publication venue or citation count, others focused on cues from within the content of the paper. For the design of Collage, we focus on allowing users to assess multiple different models for extracting the types of content that they would be interested in; we also design our tool with extensibility to new models as a primary concern.
@@
Our participants noted that they regularly work with documents across a wide time range. Several participants noted that the work that they relied on most frequently were technical reports from the 1950s to the 1970s. These reports are now digitized, but are otherwise highly variable in their accessibility to modern processing tools: The OCR used when digitizing them can be inaccurate, they often contain noise in the scanned images, and layouts are less standardized. We therefore aim to provide an interface that allows users to inspect intermediate stages of processing, to better understand where a model may have failed, and what subsequent development should target next.
@@ 3.2 Modeling and Software Interfaces
To facilitate the easy implementation of information extraction tools (in the form of PaperMage Predictors on the PDF representation), we define common interfaces that simplify the process of adding additional processing to a document’s content. These interfaces standardize three kinds of annotation on PDF content, abstracting away the need to deal with extracting PDF content or rendering it visually on the PDF, requiring users to implement only a few simple functions for processing; all models currently in Collage are implementations of these interfaces. Each of these interfaces additionally requires users to declare an identifier for the predictor so that it can be visualized in the frontend. We describe the interfaces, the requirements for implementation, and current implementations below. All interfaces are defined in the papermage_components/interfaces package of our repository. In order to add a new custom processor, users must define a class that extends one of the interfaces specified below, and then register their predictor in the local_model_config.py module.
@@
This interface is intended for any model that produces annotations of spans in text, i.e. most “classical” NER or event extraction models. Users are required to extend the TokenClassificationPredictorABC class and override the tag_entities_in_batch method, which takes a list of strings to tag, and produces a list of lists of tagged entities per-sentence. Tagged entities are expected to have the start and end character offsets, and the interface’s code automatically handles mapping indices from the sentence level to the document level.
@@
Given the prominence of large language model-based approaches, this interface is designed to allow for text-to-text prediction. Users are required to extend the TextGenerationPredictorABC class, and to implement the generate_from_entity_text() method, which takes and returns a string. This basic setup allows users to e.g. prompt an LLM and display the raw response. A popular prompting method, however, is to request structured data e.g. in the form of JSON. To accommodate this, and to allow for aggregating LLM predictions into a table, users can also implement the postprocess_text_to_dict() method. The default implementation of this method attempts to deserialize the entirety of the LLM response into a dictionary, but users can implement more specific logic if needed.
@@
Given the focus on tables and charts that many of our interview participants discussed, and the fact that table parsing is an active research area, we additionally provide an interface for models that parse images, the ImagePredictorABC in order to handle multimodal processing, including tables. This interface allows users two options of method to override: In cases where only image inputs are needed (e.g. if a table extractor performs its own OCR), the process_image() method; in cases where the method is inherently multimodal, implementors can instead override the process_entity() method, which allows them full access to PaperMage’s multimodal Entity representation. This interface requires implementors to return at least one of three types of data: a raw string representation, which we view as useful for e.g. image captioning tasks; a tabular dictionary representation, for the case of table parsing; or a list of bounding boxes, in the case of models that segment images. Implementations of this interface are free to return more than one type of output; all of them will be rendered in the frontend.
@@ 3.3 Visualization Frontend
We present the results of the PDF processing in an interactive tool built using Streamlit [^4] https://streamlit.io that allows the user to upload a PDF, define a processing pipeline, and inspect the results of that processing pipeline at each stage. More concretely, after the paper is uploaded and processed, we present the results of the pipeline in three views, in decreasing order of abstraction from the paper. The intention of this is to first show the user the potential output of their chosen pipeline for a given paper, then allow them to inspect each step of the pipeline that led to that final output. Each view is described in more detail below, and has a screenshot in Appendix A.
@@
4 Evaluation
We evaluate our system based on two measures: the degree to which it addresses the concerns shared in our interviews, and how it can be situated among existing related tools.
4.1 Addressing Needs from Interviews
@@
4.2 Comparison with Related Work
@@ 5 Conclusion
In this work, we present Collage, a system designed to facilitate rapid prototyping of mixed modality information extraction on PDF content. We focus on a case study in the materials science domain, that allows materials scientists to evaluate models for their ability to assist in literature review. We intend for this work to be a platform on which to evaluate further modeling work in this area.
@@
Our interview study was evaluated and approved by the Carnegie Mellon University Institutional Review Board as STUDY2023_00000431.

### v2

# Collage: Decomposable Rapid Prototyping for Co-Designed Information Extraction on Scientific PDFs

@@ Abstract
Recent years in NLP have seen the continued development of domain-specific information extraction tools for scientific documents, alongside the release of increasingly multimodal pretrained language models. While applying and evaluating these new, general-purpose language model systems in specialized domains has never been easier, it remains difficult to compare them with models developed specifically for those domains, which tend to accept a narrower range of input formats, and are difficult to evaluate in the context of the original documents. Meanwhile, the general-purpose systems are often black-box and give little insight into preprocessing (like conversion to plain text or markdown) that can have significant downstream impact on their results.
In this work, we present Collage, a tool intended to facilitate the co-design of information extraction systems on scientific PDFs between NLP developers and scientists by facilitating the rapid prototyping, visualization, and comparison of different information extraction models on the content of scientific PDFs. For scientists, Collage provides side-by-side visualization and comparison of multiple models of different input modalities in the context of the PDF content they are applied to; for developers, Collage allows the rapid deployment of new models by abstracting away PDF preprocessing and visualization into easily extensible software interfaces. We also enable both developers and scientists to inspect, debug, and better understand modeling pipelines by providing granular views of intermediate states of processing. We demonstrate our system in the context of information extraction to assist with literature review in materials science.
Collage: Decomposable Rapid Prototyping for Co-Designed Information Extraction on Scientific PDFs
@@ 1 Introduction
This paper presents Collage, a tool that facilitates the rapid prototyping, visualization, and comparison, of multiple models across modalities on the contents of scientific PDF documents. Collage was designed to address the interface between developers of NLP-based tools for scientific documents and the scientists who are the intended users of those tools. To address scientists’ needs, we ground our design in a series of interviews with domain experts in multiple fields, with a particular focus on materials science. Further, in cases where model results may not meet scientists’ or developers’ expectations, we visualize the intermediate representation at each step, giving the user a granular view of the modeling pipeline, allowing shared debugging processes between developers and users. Collage is domain-agnostic, and can visualize any model that conforms to one of its three interfaces - for token classification models, text generation models, and image/text multimodal models. We provide implementations of these interfaces that allow the use of any HuggingFace token classifier, multiple LLMs, and several additional models without requiring users to write any code. All of the interfaces are easily implemented, and we provide instructions and reference implementations in our repository 111github.com/gsireesh/ht-max.
@@ 2 Motivation
One of the most prominent themes to emerge in our interviews is the variety of focuses that scientists, even in very closely related subfields, can have when reading a paper and evaluating it for relevance to their purpose. While many participants focused on paper metadata, such as the reputation of the publication venue or citation count, others focused on cues from within the content of the paper. For the design of Collage, we focus on accelerating co-design of models that address specific information extraction needs on paper content, by reducing the burden of deploying new models on PDF content, and providing a shared, user-friendly view of the results upon which scientists and developers can base subsequent efforts.
@@
Our participants noted that they regularly work with documents across a wide time range. Several participants noted that the work that they relied on most frequently were technical reports from the 1950s to the 1970s. These reports are now digitized, but are otherwise highly variable in their accessibility to modern processing tools: The OCR used when digitizing them can be inaccurate, they often contain noise in the scanned images, and layouts are less standardized. This can lead to confusion on whether issues with performance are the fault of models themselves, or preprocesing choices that cause that degraded performance. We therefore aim to provide an interface that allows users to inspect intermediate stages of processing, to better understand where a model may have failed, and what subsequent development should target next: whether better performing models, or better preprocessing.
@@ 3.2 Modeling and Software Interfaces
To facilitate the easy implementation of new information extraction tools, we define common interfaces that simplify the process of adding additional processing to a document’s content. These interfaces standardize three kinds of annotation on PDF content, allowing users convenient access to the PDF’s content as images or strings (though they can access the PaperMage representation) and automatically handling visualization in several supported formats. This requires users to implement only a few simple functions in the modalities their models already use. All models currently in Collage are implementations of these interfaces. We describe the interfaces, the requirements for implementation, and current implementations below. All interfaces are defined in the papermage_components/interfaces package of our repository. In order to add a new custom processor, users must define a class that extends one of the interfaces specified below, and then register their predictor in the local_model_config.py module.
@@
This interface is intended for any model that produces annotations of spans in text, i.e. most “classical” NER or event extraction models. Users are required to extend the TokenClassificationPredictorABC class and override the tag_entities_in_batch method, which takes a list of strings to tag, and produces a list of lists of tagged entities per-sentence. Tagged entities are expected to have the start and end character offsets, and the interface’s code automatically handles mapping indices from the sentence level to the document level, and visualizing annotated text using the displaCy visualizer  [^2] https://demos.explosion.ai/displacy-ent.
@@
Given the prominence of large language model-based approaches, this interface is designed to allow for text-to-text prediction. Users are required to extend the TextGenerationPredictorABC class, and to implement the generate_from_entity_text() method, which takes and returns a string. This basic setup allows users to e.g. prompt an LLM and display the raw response. A popular prompting method, however, is to request structured data e.g. in the form of JSON. To accommodate this, and to allow for aggregating LLM predictions into a table, users can also implement the postprocess_text_to_dict() method. The default implementation of this method attempts to deserialize the entirety of the LLM response into a dictionary, but users can implement custom logic.
@@
Given the focus on tables and charts that many of our interview participants discussed, and the fact that table parsing is an active research area, we additionally provide an interface for models that parse images, the ImagePredictorABC in order to handle multimodal processing, including tables. This interface allows users two options of method to override: In cases where only image inputs are needed (e.g. if a table extractor performs its own OCR), the process_image() method; in cases where the method is inherently multimodal, implementors can instead override the process_entity() method, which allows them full access to PaperMage’s multimodal Entity representation. This interface requires implementors to return at least one of three types of data: a raw string representation, which we view as useful for e.g. image captioning tasks; a tabular dictionary representation, for the case of table parsing; or a list of bounding boxes, in the case of models that segment images. Implementations of this interface are free to return more than one type of output; all of them will be visualized in the frontend.
@@ 3.3 Visualization Frontend
We present the results of the PDF processing in an interactive tool built using Streamlit [^6] https://streamlit.io that allows the user – whether scientist or developer – to upload a PDF, define a processing pipeline, and inspect the results of that processing pipeline at each stage. More concretely, after the paper is uploaded and processed, we present the results of the pipeline in three views, in decreasing order of abstraction from the paper. The intention of this is to first show the user the potential output of their chosen pipeline for a given paper, then allow them to inspect each step of the pipeline that led to that final output. Each view is described in more detail below, and has a screenshot in Appendix A.
@@
4 Addressing Needs from Interviews
@@
5 Co-design with Collage
In this section, we walk through an example of how Collage might facilitate the development of an information extraction pipeline for a materials scientist. In this scenario, Bob, a materials scientist, wishes to extract the synthesis parameters of a class of materials called zeolites from a dataset of PDFs from the 1980s to the 2010s. Papers discussing Zeolite synthesis often report parameters both in the text of the paper as well as in tables, so multimodal extraction is crucial. He has worked with Alice, an NLP developer, before but they have not yet collaborated on this project.
Evaluating off-the-shelf models.
Bob begins in Collage by trying to see if there is an existing model that already works for his case. Using the HuggingFace model selector in the upload paper view, he searches for tagging models, but only finds models trained on general scientific or biomedical text, not materials. He is, however, able to write a prompt for an LLM model to extract this information, and he adds predictors that call out to two popular commercial models to extract the information that he’s interested in. He uploads a recent paper that he’s been reading, and waits for Collage to process it.
Finding modeling opportunities.
Once Collage has processed the paper, Bob heads to the summary view, and compares the results from the two commercial models. He’s able to view the parameters that they extracted, filtered by section, to develop an understanding of what heuristics might get him the information he wants: parameters identified in the related works section, for example, are frequently irrelevant to his search. In the summary view, he’s also able to see the tables that Collage has identified and parsed with the TableTransformer and MathPix models, along with their labels and captions, and the tagged bounding boxes for the table cells.
To make sure those annotations are reasonable, he heads to the annotations view, where he can visualize the extracted information side-by-side with the original PDF content, and compare the annotations from his two LLMs. He’s also able to check whether the table detection model has predicted sensible bounding boxes that both don’t exclude content like table footnotes, but also don’t include irrelevant, non-table content. He notes that while the table parsing from both models is reasonable, the paper he’s reading reports values in ratios that may not be comparable across papers. To have a single pipeline that produces normalized results, he’d like to use a multimodal LLM, but in Collage currently, LLMs can only be applied to text. He decides to get in touch with Alice, to see if she can develop an LLM-based table information extraction model.
Prototype model development.
Alice begins work on a table information extraction tool, but there are a lot of possible options to evaluate: should she use a multimodal model and process the table in image format? Should she linearize the table into text, and have a text-only LLM work with it? In Collage, both options involve little more than implementing the LLM call, so it’s easy to do both and then compare. For the multimodal case, Alice extends the image predictor interface, which allows her to receive as input the cropped image of any element on the page and pass that to an LLM; for the text-only case, Alice can easily access the underlying document representation use the already identified and parsed tables (which are in a DataFrame-compatible format) and convert them into markdown for her linearization. She is able to return a dictionary in the same schema for both predictors, which will automatically be visualized in the frontend as a Pandas dataframe. She commits her code, registers the predictors, and asks Bob to take a look in the Collage interface.
In-context evaluation.
Bob then re-processes his paper through Collage, making sure to check the boxes for Alice’s new table parsing predictors. In the summary view, he’s able to compare the predicted, normalized tables to the original PDF, to verify that the models are performing the normalization correctly. He then picks the better performing model, and asks Alice to create a pipeline that can process his entire dataset. Alice is able to take the predictor, add it to the PaperMage recipe that underlies Collage, and run it over Bob’s set of PDF documents, adding a step to export the parsed tables that Bob saw in the Collage interface.
Diagnosing errors.
Bob looks through the parsed tables from processing all of the PDFs, and notes that for the older PDFs, the parsed content doesn’t look right. He’d like to diagnose the problem. Because the processing that Alice and Bob run on these documents is the same as that underlying Collage, the results can be visualized in the tool, even if they were not directly processed through it. Bob loads the representation of the parsed older document, and is able to view the results from the model that didn’t look right. While the bounding boxes for the table look correct in the annotations view, he’s also able to see in the inspection view that the text detected within the table has not been correctly OCR’d. He can now contact Alice to see if there’s a fix for that problem, but in the meantime, he can examine the visualizations for his PDFs to understand how the publication year might affect whether the deployed suite of models can correctly extract and normalize information, and what the cutoff year might be for the results to be trustworthy.
In this case, Collage enables Bob to self-serve cutting-edge NLP for his own use case, requiring that he involve Alice only when Collage’s functionality needs extension. When that happens, Bob and Alice can both see results in the same interface, and can discuss errors and how to prioritize new work. When Alice develops new predictors to address Bob’s needs, she is required to do no PDF processing or visualization, which are built into the tool, and Bob can evaluate and compare the results of these new predictors in the same interface he’s been using the whole time. For debugging, both Bob and Alice have access to the same representation and visualization as a shared source of truth, and collaborate to involve both NLP and subject matter expertise in how to fix the problem. Collage can accelerate the process of collaboration between NLP developers and scientists, allowing for co-design and rapid prototyping with a shared representation.
6 Related Work
@@ 7 Conclusion
In this work, we present Collage, a system designed to facilitate co-design and rapid prototyping of mixed modality information extraction on PDF content between scientists and NLP developers. We focus on a case study in the materials science domain, that allows materials scientists to evaluate models for their ability to assist in literature review. We intend for this work to be a platform on which to evaluate further modeling work in this area.
@@
Our interview protocol was evaluated and approved by the Carnegie Mellon University Institutional Review Board as STUDY2023_00000431.